$ cost and Ev
I think tucson would like to see the exact EV of the GN blackjack tournament. To do so, other than the correct formulas provided by other posters in this thread, we should use concrete format with specific numbers of players. Also, we need to clarify if one attempts to just pay the original entry fees or rebuy and superrebuy, as well.
From the link Colin provided it shows that there are 5 players per table in the semifinals.
Let’s make a model, which may be not as exact as the one of the coming tournaments but should serve as an example.
We plan on all rebuys, 4 out of 5 drawn names are present (in many tourneys it is much lower).
The format could be as follow:
1st Round
162 players – 27 Tables (all tables have 6 players, two advance)
54 players advance, 108 don’t advance
Rebuy
90-100 people rebuy, 18 tables, 6 or 5 per table (average 5.1), 36 advance
Wild Card to 2nd round: 18
total advancement to 2nd round 108
2nd round
108 players, 18 tables of 6, 36 advance
Superrebuy 66-72 players, 12 tables, mostly 6 per table, 24 advance
WC to 3rd round: 6 names
total advancement to quarterfinals around 66
Quarterfinals
66 players, 11 tables, 22 advance
WC to semis: 3 names
total in semifinals 25 players
Semifinal
25 players, 5 tables of 5, one per table advances to the final
plus 1 Wild card
We can calculate the chances of advancing for the average skill player:
to 2nd round:
1st play 33.33%
Rebuy (66.67x 39.2%) = 26.14%
WC 40.5% x (18/58) = 12.6%
Total chance 72%
to quarters:
1st play 72% x 33.33% = 24%
superrebuy (72%x.66.66%) x 33.33 = 16%
WC 60% x 6/82 = 4.4%
Total chance 44.4%
to semifinals:
44.4% x 33.33% = 14.8%
WC 85.2% x 3/112 = 2.3%
Total chance 17.1%
to final:
17.1% x 20% = 3.4%
WC 96.6% x 1/120 = 0.8%
Total chance to make the final 4.2%
As you can see making finals happens once in about 24 tournaments, which is slightly better than 6/162 (once in 27) because not all players rebuys and show for the wild cards.
The total expected value is $696 ($612 from the final table prizes).
The expected cost because of all possible rebuys is $1,073.
So, for people who pay all the entries it is 35% disadvantage.
However, I don’t know a single person who thinks that his skills are below the average player. For people who are better than average it is quite another story.
S. Yama